OpenStudy (anonymous):

HELP PLEASE NEED TO STUDY FOR MIDTERMS ' 10x + 4 and 11x - 4 solve for x . I'm about to draw it for you also

OpenStudy (anonymous):



OpenStudy (anonymous):

@Luis_Rivera I meant 10x + 4

OpenStudy (anonymous):

what ? how did you get that . don't you have to put them equal to eachother ... like 11x - 4=10x + 4 then solve .... @Luis_Rivera

Directrix (directrix):

@Cee69 If these expressions are measures of alternate interior angles formed by a transversal which cuts two lines AND if the task is to find the value of x which would make the two angles equal in measure, then proceed with what you wrote: --> 11x - 4=10x + 4 then solve

OpenStudy (anonymous):

startt off with 11 - 10= 1x then -4 - 4= -8 so now u have x=-8

Directrix (directrix):

How do you know that 11x - 4 + 10x + 4 =180 is not the correct equation for this problem? @star95 if 11x - 4 =10x + 4 is the correct one, check your work. I think -8 is not correct. The error lies here: then -4 - 4= -8 (copied from your work)

OpenStudy (anonymous):

idk sometimes i get things mixed up... last night i was stressed, didnt mean to get it wrong

Directrix (directrix):

@star95 I understand. If you want to work on the problem, look at the given instructions, post them, and we'll find the answer. Just guessing, but I think the instructions will state something about finding the value of x which makes the two given lines parallel.